# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

### v2

@@ 4 Achieving Linear Speedup for Push-Pull
For simplicity, we denote $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$. Using the gradient tracking recursion (8b), we discover the following relationship about extra error terms: for any integer $k\geq 0,m\geq 1$:
$\displaystyle\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (11)
To understand this equation, we propose the following important supporting lemma:
@@
For any $l\geq 0$ and any matrices $M^{(0)},M^{(1)},\ldots,M^{(l)}\in\mathbb{R}^{n\times n}$. For any $k\geq 0$, If $H=\sum_{i=0}^{l}M^{(i)}{\mathbf{g}}^{(k+i)}$ is a linear combination of stochastic gradients, then
Proof.
This is a natural corollary of Assumption 3. ∎
Now we turn to the noise estimation for $\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$. A common analysis using Cauchy’s inequality typically provides
$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\Delta_{y}^{(k+i)}\|_{F}^{2}]\lesssim m^{2}\sigma^{2}.$ (12)
Here, we use $\lesssim$ to absorb matrix-related constants and focus on $\sigma^{2}$-related terms. Nevertheless, analysis using Equation (11) and Lemma 3 can be significantly better. Proposition 2 ensures that $\textstyle\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}+\|\sum_{j=1}^{m-1}B^{m-j-1}-B_{\infty}\|_{2}^{2}$ is finite. Therefore, applying Lemma 3 to Equation (11) tells us
$\displaystyle\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]\lesssim\mathbb{E}_{k}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+\sigma^{2}$ (13)
$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\lesssim\sigma^{2}+\alpha^{2}L^{2}m^{2}\sigma^{2}.$
Estimate from (13) sufficiently exploits the structure of gradient tracking, thus exhibiting ${\mathcal{O}}(m^{2})$ improvement compared to a naive estimate (12). To effectively leverage the result of  (13), we introduce an $m$-step descent lemma. It comes from applying the standard descent lemma to the following recursion:
$\hat{x}^{(k+m)}=\hat{x}^{(k)}-c\alpha\sum_{i=0}^{m-1}\bar{g}^{(k+i)}-\alpha\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}.$ (14)
Lemma 4 (m-step Descent Lemma).
Suppose Assumption 34 and Proposition 1 hold. For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$. When $\alpha\leq\frac{1}{6cmL}$, the following inequality holds for any integers $k\geq 0,m\geq 6c^{-2}s_{B}^{2}$:
$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (15)
$\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}$
$\displaystyle\quad+\frac{c\alpha L^{2}}{2}\Delta_{1}^{(k)}+\frac{c\alpha L^{2}}{2}\Delta_{2}^{(k)}+\frac{3\alpha s_{B}^{2}}{cm}\|\Delta_{y}^{(k)}\|_{F}^{2},$
where
$\displaystyle\Delta_{1}^{(k)}=\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}=\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
The remaining errors are estimated in the next lemma.
Lemma 5 (Remaining Errors).
@@
$\displaystyle\quad\frac{c\alpha L^{2}}{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{3\alpha s_{B}^{2}}{cm}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq\frac{\alpha c}{4}\sum_{t=0}^{mK-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(t)}\|^{2}]+\frac{\alpha cm}{4}(\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}})\sigma^{2}+C_{\Delta,1}\frac{c\alpha L\Delta}{4m}.$
where we define $[m^{*}]:=\{0,m,\ldots,(K-1)m\}$, and $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants. $s_{B^{m}}$ is the measure defined for $B^{m}$, see (3). Generally, we have $s_{B^{m}}\leq s_{B}$, and $s_{B^{m}}$ converges to $\|I-B_{\infty}\|_{2}$ for large $m$.
We are ready to state the main result of this paper.
@@
where $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants. Furthermore, if we choose
@@
See Appendix B.3
Comparison with existing analyses. Theorem 1 provides the first finite-time non-convex convergence result for stochastic Push-Pull, while previous analyses focus either on the strongly convex case [46, 40] or on variants of Push-Pull relying on specific topologies [5]. The first term $\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ explicitly shows that Push-Pull enjoys a linear speedup on arbitrary topologies satisfying Proposition 2.
Impact of $m$: In our analysis, we consider $m$-step gradient descent. The parameter $m$ effectively controls the terms related to $\sigma^{2}$ from additional errors, as illustrated by the $\mathcal{O}(m^{-2}\sigma^{2})$ term in (16). $m$ cannot be arbitrarily large because larger values of $m$ restrict the range of the learning rate, resulting in slower convergence and longer transient times in the analysis. In summary, the parameter $m$ allows for a tradeoff between noise level and transient time.
Transient time. Our theorem is able to provide a coarse estimate for the transient time, which is
$\max\{n^{2}c^{-4}s_{B}^{8},n^{7}c^{-4}s_{A}^{2}s_{B}^{3},nc^{-4}s_{B}^{4}\}.$
The order of $n$ is overestimated because we frequently use $\|\pi_{A}\|\leq 1$ and $\|\pi_{B}\|\leq 1$, leading to a loss of $n$ when $A$ or $B$ is doubly-stochastic.
@@ 5 Experiment
In this section, we empirically investigate the linear speedup properties of the Push-Pull algorithm across diverse network topologies. For the stochastic gradient oracle, we ensure that each node operates on a local dataset, computing gradients from randomly sampled data points at each iteration. The experimental validation proceeds in two phases: first, we examine a nonconvex regression problem using synthetic data, followed by training a fully connected neural network on the heterogeneously distributed MNIST dataset.
To further assess practical performance on established benchmarks, we extend our evaluation to ResNet-18 training on CIFAR-10. The distributed training experiments are conducted across multiple nodes ($n>1$) using three distinct network topologies: directed exponential graphs, grid graphs, and undirected random graphs. All results are systematically compared against the single-node baseline ($n=1$) to quantify the speedup achieved through distributed optimization.
Complete implementation specifications are provided in Appendix C. The source code and experimental scripts are publicly available in our GitHub repository.
@@ 5.1 Nonconvex Logistic Regression on Synthetic Dataset
In the first experiment, the loss function at each node is a logistic regression function with nonconvex regularization terms (see (44) in Appendix C). We studied six different network topologies: the exponential, grid, and ring graphs are directed, while the random, geometric, and nearest neighbor graphs are undirected. The mixing matrices $A$ and $B$ are generated by a special procedure (see Appendix C.1), which produces row-stochastic or column-stochastic matrices, but not doubly-stochastic ones.
We can observe clear linear speedup curves in Figure 2. For each topology, the noise level in gradients decreases proportionally to the square root of the number of nodes after the same number of communication rounds. All learning rates are set to $lr=0.005$.
Figure 2: Linear speedup performance over various topologies on synthetic data, the learning rates are all 0.0050.005 0.0050.005.
@@ 5.2 Training Neural Network for MNIST Classification
In the second experiment, we trained a three-layer fully connected neural network for MNIST classification. We used the same topologies as before: directed exponential, grid, and ring graphs, and undirected random, geometric, and nearest neighbor graphs. Data is distributed heterogeneously using a Directlet distribution, see Appendix C.3. Figure 3 shows that Push-Pull is able to achieve linear speedup on both directed and undirected topologies.
@@ 5.3 Training Neural Network for CIFAR10 Image Classification
In the third experiment, we trained Resnet18 for CIFAR10 classification across different numbers of nodes and various topologies. We used three topologies mentioned before: directed exponential, grid graphs, and undirected random graphs. Here $n=1$ means training a single Resnet18 on a single-node without distributed optimization.
Figure 4 demonstrates the performance of our Push-Pull method. We observe that Push-Pull’s accuracy is initially lower compared to single-node optimization, which can be attributed to the overhead of the influence of topology. Nevertheless, after a sufficient number of epochs, Push-Pull achieves superior accuracy, leveraging its inherent speed advantages.
Figure 4: The test accuracy across different numbers of nodes and various topologies. All experiments were conducted using a learning rate of 0.0050.005 0.0050.005. The lower three figures specifically depict the accuracy after 100100 100100 epochs.
@@ 6 Conclusion
Appendix A Rolling-Sum Lemma
@@
[…]
@@
[…]
@@ B.1 Proof of Lemma 4
For simplicity, we define $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$ and $\overline{\nabla f}^{(k)}:=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x^{(k)}_{i})$. According to Assumption 3, we have
@@
By taking expectation conditioned on ${\mathcal{F}}_{k}$, we obtain that
@@
Plug these estimations into (24), we obtain that
@@
$\displaystyle\quad+\mathbb{E}_{k}[\|\left(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\right)({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
@@
The inner product can produce a similar term:
$\mathbb{E}_{k}[\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}]=\mathbb{E}_{k}[H_{m,k}].$
Using the fact that $\hat{x}^{(k)}$ is known under filtration ${\mathcal{F}}_{k}$, we have
$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]=-\alpha\left\langle\mathbb{E}_{k}[H_{m,k}],\nabla f(\hat{x}^{(k)})\right\rangle$ (28)
@@
[…]
$\displaystyle\quad+\frac{2\alpha}{cm}\|\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)}\|_{F}^{2}+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}$
@@
$\displaystyle+\frac{3\alpha}{cm}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
We can further apply Cauchy’s inequality on the last term in (31), i.e.
@@
Finally, when $m\geq 6c^{-2}s_{B}$, we have $\frac{3\alpha s_{B}L^{2}}{cm}\leq\frac{c\alpha L^{2}}{2}$, and (31) can be further simplified to
@@
which finishes the proof.

### v3

@@ 5 Achieving Linear Speedup for Push-Pull
In this section, we formalize the aforementioned arguments through supporting lemmas and rigorously establish the linear speedup property for Push-Pull.
5.1 Supporting lemmas
Lemma 3 (m-step Descent Lemma).
Suppose Assumption 34 and Proposition 1 hold. For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$. When $\alpha\leq\frac{1}{6cmL}$, the following inequality holds for integers $k\geq 0,m\geq 6c^{-2}s_{B}^{2}$ (Proof is in Appendix B.1):
$\displaystyle\quad\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{1}{m}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (22)
$\displaystyle\leq\frac{4(f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})])}{\alpha cm}+\frac{8\alpha cL}{n}\sigma^{2}+\frac{8\alpha Lns_{B}^{2}}{cm}\sigma^{2}$ (23)
$\displaystyle\quad+\frac{2L^{2}}{m}\Delta_{1}^{(k)}+\frac{2L^{2}}{m}\Delta_{2}^{(k)}+\frac{12s_{B}^{2}}{c^{2}m^{2}}\|\Delta_{y}^{(k)}\|_{F}^{2},$ (24)
where $\overline{\nabla f}^{(k+i)}:=n^{-1}\mathbf{1}^{\top}\nabla f({\mathbf{x}}^{(k+i)})$ denotes the average of the local deterministic gradients evaluated at ${\mathbf{x}}^{(k+i)}$, and
$\displaystyle\Delta_{1}^{(k)}:=\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}:=\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
In (23), when $m\geq c^{-2}n^{2}s_{B}^{2}$, the noise terms associated with $\sigma^{2}$ are dominated solely by the linear speedup term $\tfrac{8\alpha cL}{n}\sigma^{2}$, which is consistent with our analysis in (21). The remaining error terms appearing in (24) are estimated in the next lemma.
@@ 5.1 Supporting lemmas
$\displaystyle\quad\frac{2L^{2}}{m}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{12s_{B}^{2}}{c^{2}m^{2}}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq\frac{1}{m}\sum_{t=0}^{T-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(t)}\|^{2}]+(\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}})\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{4m^{2}}.$
Here, $C_{m}$ and $C_{\Delta,1}$ denote matrix-related constants (see (B.2)). The quantity $s_{B^{m}}$ represents the matrix measure of $B^{m}$, as defined in (5). In general, we have $s_{B^{m}}\leq s_{B}$, and moreover, $s_{B^{m}}$ converges to $\lVert I-B_{\infty}\rVert_{2}^{2}$ as $m\to\infty$ (Proof is in Appendix B.2).
With the above lemmas, we are ready to state the main result of this paper.
@@
where $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants defined in (B.2) and (53). Furthermore, if we choose
@@
See Appendix B.3. ∎
Comparison with existing analyses. Theorem 1 provides the first convergence guarantee for stochastic Push-Pull on nonconvex problems. In contrast, prior analyses focus either on the strongly convex setting [42, 25] or on Push-Pull variants restricted to specific network topologies [7]. The leading term $\tfrac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ demonstrates a linear speedup with respect to the number of nodes $n$ for any topology satisfying Proposition 2. Consequently, Push-Pull requires $\mathcal{O}(1/(n\epsilon^{2}))$ iterations to achieve an $\epsilon$-accurate solution, with the complexity decreasing linearly as $n$ increases.
Impact of $m$: In our analysis, we consider $m$-step gradient descent, where the parameter $m$ governs the noise-related terms, as reflected by the $\mathcal{O}(m^{-2}\sigma^{2})$ term in (25). However, $m$ cannot be taken arbitrarily large: increasing $m$ reduces the admissible range of the learning rate, thereby slowing convergence in the analysis. Overall, $m$ balances noise reduction against convergence rate. With an appropriate choice of $m$, one can suppress noise, achieve linear speedup, and still guarantee $\mathcal{O}(1/\sqrt{nT})$.
@@ 6 Experiment
In this section, we empirically examine the linear speedup properties of the Push-Pull algorithm across a variety of network topologies. Each node in the stochastic gradient oracle operates on a local dataset, computing gradients from randomly sampled data points at every iteration. The experimental study consists of two phases: (i) nonconvex regression on synthetic data, and (ii) training a fully connected neural network on the heterogeneously distributed MNIST dataset.
To further evaluate practical performance on standard benchmarks, we extend our experiments to training ResNet-18 on CIFAR-10. The distributed training is performed across multiple nodes ($n>1$) under different network topologies, including directed exponential graphs, grid graphs, and random graphs. For reference, all results are systematically compared against the single-node baseline ($n=1$) to quantify the speedup obtained through distributed optimization.
Complete implementation details are provided in Appendix C, and the source code with experimental scripts is publicly available at our GitHub repository.
@@ 6.1 Nonconvex Logistic Regression on Synthetic Dataset
In the first experiment, the loss function at each node is logistic regression with nonconvex regularization terms. We evaluate six network topologies: exponential, grid, and ring graphs (directed), and random, geometric, and nearest-neighbor graphs (undirected). The mixing matrices $A$ and $B$ are constructed using a specialized procedure, which yields row-stochastic or column-stochastic matrices, but not doubly stochastic ones. Details are provided in Appendix C.1 and Appendix C.2.
As shown in Figure 2, the results exhibit clear linear speedup. Across all topologies, the gradient noise level decreases proportionally to the square root of the number of nodes, given the same number of communication rounds. The learning rate is fixed at $lr=0.005$ for all experiments.
Figure 2: Linear speedup performance over various topologies on synthetic data, the learning rates are all $0.005$.
@@ 6.2 Training Neural Network for MNIST Classification
In the second experiment, we trained a three-layer fully connected neural network for MNIST classification. The same network topologies as in the synthetic dataset experiment were used. The MNIST data are distributed heterogeneously according to a Dirichlet distribution (see Appendix C.3). As shown in Figure 3, Push-Pull achieves linear speedup on both directed and undirected topologies.
@@ 6.3 Training Neural Network for CIFAR10 Image Classification
In the third experiment, we trained ResNet-18 for CIFAR-10 classification across varying numbers of nodes and network topologies. We considered three representative topologies: directed exponential graphs, grid graphs, and undirected random graphs. For reference, $n=1$ corresponds to training ResNet-18 on a single node without distributed optimization.
Figure 4 illustrates the performance of the Push-Pull method. We observe that its accuracy is initially lower than that of single-node training, due to the overhead introduced by network topology. However, after a sufficient number of epochs, Push-Pull surpasses the single-node baseline, benefiting from its inherent speedup advantages.
Figure 4: The test accuracy across different numbers of nodes and various topologies. All experiments were conducted using a learning rate of $0.005$. The lower three figures specifically depict the accuracy after $100$ epochs.
@@ 7 Conclusion
Appendix A Linear Algebra Inequalities
@@
$\displaystyle\leq$ $\displaystyle\sum_{i=0}^{\min\{\lfloor p\rfloor,k\}}M_{A}+\sum_{i=\min\{\lfloor p\rfloor,k\}+1}^{k}M_{A}\beta_{A}^{i-p}$
$\displaystyle\leq$ $\displaystyle M_{A}\cdot(1+\min\{\lfloor p\rfloor,k\})+M_{A}\cdot\frac{1}{1-\beta_{A}}\beta_{A}^{\min\{\lfloor p\rfloor,k\}+1-p}.$
@@
$\displaystyle\leq$ $\displaystyle\sum_{i=0}^{k}a_{k-i}\|a_{k-i}^{-1}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq\sum_{i=0}^{k}a_{k-i}^{-1}\|A^{k-i}-A_{\infty}\|_{2}^{2}\|\Delta^{(i)}\|_{F}^{2}.$
@@
Lemma 6.
Given $l\geq 0$ and matrices $M^{(0)},M^{(1)},\ldots,M^{(l)}\in\mathbb{R}^{n\times n}$. For any $k\geq 0$, If $H=\sum_{i=0}^{l}M^{(i)}{\mathbf{g}}^{(k+i)}$ is a linear combination of stochastic gradients, then
Proof.
This is a direct corollary from the unbiasedness and linearly independent gradients in Assumption 3. ∎
@@ B.1 Proof of Lemma 3
For simplicity, we define $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$ and $\overline{\nabla f}^{(k)}:=n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})$. According to Assumption 3, we have
@@
Take expectation conditioned on ${\mathcal{F}}_{k}$ and we obtain that
@@
Plug the estimation for the inner product into (33), we obtain that
@@
$\displaystyle\quad+\mathbb{E}_{k}[\|(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
@@
Similarly, using the fact that $\hat{x}^{(k)}$ is known under filtration ${\mathcal{F}}_{k}$, we have
$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (37)
@@
$\displaystyle\overset{\text{Cauchy}}{\leq}\frac{2\alpha}{cm}\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|^{2}]$ (38)
$\displaystyle\quad+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{2\alpha}{cm}\|\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)}\|_{F}^{2}+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}$
@@
$\displaystyle+\frac{3\alpha}{cm}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}].$
We can further apply Cauchy’s inequality on the last term in (41), i.e.,
@@
Finally, when $m\geq 6c^{-2}s_{B}$, we have $\frac{3\alpha s_{B}L^{2}}{cm}\leq\frac{c\alpha L^{2}}{2}$, and (41) can be simplified to
@@
By dividing $\frac{\alpha cm}{4}$ on both sides of this inequality, we finish the proof.